3. Consider two stocks, A and B, with their expected returns and standard deviations, as follows: Stock Expected return Stand letter b please
Answer #1

A.

Weight A=1/2=       0.5  
          
Standard deviation of A (σA)=       10%  
          
weight B=1/2=       0.5  
Standard deviation (σB)=       8%  
Correlation between A and B=       1  
  
(σp)=   ((weigh A * σA ) ^2 + (weight B* σB )^ 2 + (2 * Weight A* Weight B*σA *σB* correlation))^(1/2)      
(((0.5*10%)^2+(0.5*8%)^2+(2*0.5*0.5*10%*8%*1)))^(1/2)          
9.00%
          
Standard deviation of portfolio is 9%          

B.

Weight A=1/2= 0.5

Standard deviation of A (σA)= 10%

weight B=1/2= 0.5

Standard deviation (σB)= 8%

Correlation between A and B= 0.5

(σp)= ((weigh A * σA ) ^2 + (weight B* σB )^ 2 + (2 * Weight A* Weight B*σA *σB* correlation))^(1/2)

(((0.5*10%)^2+(0.5*8%)^2+(2*0.5*0.5*10%*8%*0.5)))^(1/2)

7.81%

Standard deviation of portfolio is 7.81%

C.

Weight A=1/2=       0.5
      
Standard deviation of A (σA)=       10%
      
weight B=1/2=       0.5
Standard deviation (σB)=       8%
Correlation between A and B=       -0.5

(σp)=   ((weigh A * σA ) ^2 + (weight B* σB )^ 2 + (2 * Weight A* Weight B*σA *σB* correlation))^(1/2)  
(((0.5*10%)^2+(0.5*8%)^2+(2*0.5*0.5*10%*8%*-0.5)))^(1/2)      
4.58%      
      
Standard deviation of portfolio is   4.58%  
      

Please thumbs up.